Circular No. 8563

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005cv IN UGC 1359 Further to IAUC 8557, D. R. Madison and W. Li report the LOSS

discovery of an apparent supernova (mag 17.6) on unfiltered KAIT images taken on July 10.49 and 11.47 UT. SN 2005cv is located at α = 1^{h}53^{m}42^{s}.56, δ = +29^{o}56′04′′.3 (equinox 2000.0), which is 4′′.1 east and 4′′.6 north of the nucleus of UGC 1359. A KAIT image taken on Jan. 21.19 showed nothing at this position (limiting mag ∼ 19.5).

1994 XD Further to IAUC 8329, L. A. M. Benner, M. C. Nolan, S. J. Ostro, J.

D. Giorgini, J. L. Margot (now at Cornell University), and C. Magri report that Arecibo delay-Doppler images (2380 MHz, 12.6 cm), obtained on June 24–25, reveal that 1994 XD is a binary system. Preliminary estimates, based on range extents at 150-m resolution, give the primary’s diameter as ∼ 600 m and a lower bound on the secondary’s diameter as 150 m (apparent separation ∼ 600 m on June 24, with uncertainty as much as 300 m).

COMETS C/2005 K9, 2005 K10, 2005 L5, 2005 L6 (SOHO) Following are the initial available positions for additional presumed

comets found on SOHO website images (continuation to IAUC 8551). All were Kreutz sungrazers except for C/2005 K9 (Meyer group). K. Battams writes that C/2005 K9 appeared stellar, reaching mag ∼ 7.9 at 6.0R_{⊙} on May 28.189 UT. C/2005 K10 was very diffuse and had a very faint tail that was 150′′ long at 6.6R_{⊙} on May 30.481. C/2005 L5 and C/2005 L6 were both faint and diffuse with no tail (C/2005 L5 also decribed as small). C/2005 L6 was found independently by S. Hönig.

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 K9 May 28.022 4^{h}21.^{m}8 +22^{o}10′ C2 RK 2005-N60
C/2005 K10  30.439 4 30.0 +20 06 C2 KB 2005-N60
C/2005 L5  June 1.481 4 38.0 +20 21 C2 HS 2005-N60
C/2005 L6  5.772 4 54.1 +20 53 C2 TH 2005-N60

COMET P/2005 N2 (HERGENROTHER) E. J. Christensen reports that CCD images taken with the 0.5-m Up-

psala Schmidt telescope on July 5.57–5.69 UT show a moderately diffuse asymmetric coma (long axis 10′′, short axis 6′′), elongated toward p.a. 260^{o}.

2005 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green